OpenStudy (anonymous):

What is the distance between (2, 0) and (5, 8)?

OpenStudy (anonymous):

\[distance = \sqrt{(x2-x1)^2+(y2-y1)^2}\]

OpenStudy (chaise):

d = sqrt((x-x)^+(y-y)^2 d = sqrt(3^2+8^2) d = sqrt(9+64) d = sqrt(75)

OpenStudy (anonymous):

Chaise I came up with d=√(〖(5-2)〗^2+〖(8-0)〗^2 ) =√(〖(3)〗^2+〖(8)〗^2 ) =√(9+64) =√73 Did I attempt correctly?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you are right

OpenStudy (chaise):

My bad. Basic error. >.< You're right.